OpenStudy (anonymous):

How do you find the integral of tan^4(x) dx? I assume you split it into sin^4(x)/cos^4(x), but after that, how do you integrate?

OpenStudy (anonymous):

use by parts

OpenStudy (anonymous):

integration by parts

OpenStudy (anonymous):

convert tan into sec

OpenStudy (anonymous):

tan^2*(sec^2-1)

OpenStudy (anonymous):

u=tanx du=sec^2

OpenStudy (anonymous):

Problem is, I don't really know how to deal with sec. Is there any other way I can integrate by parts?

OpenStudy (anonymous):

is it mandatory to split tan into sin and cos?

OpenStudy (anonymous):

it is easier just using tan into sec

OpenStudy (anonymous):

tan^4=tan^2(sec^2-1)

OpenStudy (anonymous):

tan^2sec^2-tan^2

OpenStudy (anonymous):

integrate each one separately

OpenStudy (anonymous):

tan^2sec^2 ----> u=tan du=sec^2, so... ----> \[\int\limits_{}u ^{2}du\]

OpenStudy (anonymous):

and \[\int\limits_{}\tan ^{2}xdx=tanx-x\]

OpenStudy (anonymous):

i've just solved it for you, did you understand?

OpenStudy (anonymous):

one sec. Let me go through it :)

OpenStudy (abb0t):

Yes, split them. \(\int\limits \tan^2(x)\tan^2(x)dx \). Then, you can convert one to a trig identity (which you should recognize). Then simple u-sub.

OpenStudy (anonymous):

@SerikMB , I'm a little confused because my textbook says I should get 1/3(sinx/cosx) - 4/3 (sinx/cosx) + x + C ... I follow your process, I just don't see how they're the same.

OpenStudy (anonymous):

then you need to find it by splitting

OpenStudy (anonymous):

OK. Well thanks for walking me through the other way of doing it, though, it's good for me to know how! :)

OpenStudy (anonymous):

\[\int\limits_{}\sin ^{2}/\cos ^{2}\]

OpenStudy (anonymous):

then use integration by parts

OpenStudy (anonymous):

no use just substition

OpenStudy (anonymous):

OK. I think I can take it from there, just one question: which do I use as U, and which do I use as V'?

OpenStudy (anonymous):

u=cos du=-sin

OpenStudy (anonymous):

Ohok

OpenStudy (anonymous):

do you know why?

OpenStudy (anonymous):

because it makes it way simpler than doing it by parts?

OpenStudy (anonymous):

It makes sense that cos would be the u because then du could be -sin which simplifies everything

OpenStudy (anonymous):

ohh, sorry i messed up here. We need to use BY PARTS at final. No more messes promise)

OpenStudy (anonymous):

Wait so no substitution? Why not?

OpenStudy (anonymous):

because there is a square

OpenStudy (anonymous):

i mean 4th degree

OpenStudy (anonymous):

Right, and you can't deal with that because it won't clear sin^2(x) if you do substitution. OK, that makes sense.

OpenStudy (anonymous):

so if it's by parts, what is U and what is V'?

OpenStudy (anonymous):

wait a minute plz

OpenStudy (anonymous):

sure thing

OpenStudy (anonymous):

\[\int\limits_{}\sin ^{4}*\frac{ 1 }{ \cos ^{4} }\]

OpenStudy (anonymous):

u=sin^4 and dv=sec^2

OpenStudy (anonymous):

v=tanx and du=4sin^3cos

OpenStudy (anonymous):

wait, i again faced up with different answer

OpenStudy (anonymous):

i have to go now, but thanks for your help!! I think I'm getting it now :)